Solve the inequality: x + 14 < -11
Found 3 solutions by babygirlwalton, jim_thompson5910, lenny460:
Answer by babygirlwalton(3)
Answer by jim_thompson5910(35256)

x + 14 < -11

x + 14 -14 < -11-14

x + 0 < -25

x < -25

Answer by lenny460(1073)

Solve the inequality: x + 14 < -11

x + 14 < -11
Subtract 14 from each side of the inequality.
x + 14 - 14 < -11 - 14
Combine like terms:
14 - 14 = 0
-11 - 14 = -25
x < -25
The Answer:
x < -25



Lennox Obuong
Algebra Student
Email: obuong3@aol.com